OpenStudy (anonymous):

focus, vertex, and directrix for 16(x-9)=(y+9)^2

OpenStudy (campbell_st):

well this is in the form (y - k)^2 = 4a(x - h) its a parabola that is open to the right with the line of symmetry parallel to the x axis here is an example  so the vertex is at (h, k) so 4a = 16 a is the focal length... so find a the focus will be (h + a, k) the directrix is x = h - a hope this helps